# Leganés, Spain 22affiliationtext: Instituto de Investigación Sanitaria Gregorio Marañón Madrid, Spain

### v2

@@ G.1 Signal Reconstruction and Imputation
Figure 11 presents reconstructed and imputed sample examples, where white shading indicates observed data, gray shading denotes originally missing data, and purple shading represents synthetically induced missingness. The remaining time steps (in this case, days) are fully visible to the model. When the original signal is obscured in observed intervals, it is due to one or more model reconstructions perfectly overlapping the true signal, demonstrating accurate recovery. As shown in Figures 10(a) and 10(b) all models perform well with binary variables.
Notably, the proposed VQ-TSFM and its variants E1 and E2 exhibit strong imputation capabilities even under high proportions of missingness, as evidenced by Figures 10(d), 10(f), 10(i), and 10(g). Whether the missing data spans large temporal segments (e.g., the first three-quarters of the sample in 10(f)), appears centrally (10(i)), or is intermittently distributed (10(g)), the models consistently maintain robust representations and plausible imputations. This performance generalizes across all variable types—continuous real-valued, continuous positive, count data, and binary—highlighting the versatility of the models across different data ranges and types. The implicit model—the simplest of the three proposed variants—exhibited performance on par with or better than that of E1 and E2. This indicates that, even without explicit conditioning on the missingness mask, VQ-TSFM effectively captures and reconstructs the underlying patterns of missing data.
@@ G.2 Embedding Usage Histograms
Sleep Start (s) XO $-6.1860$ $3\times 10^{-8}$ (✗) $-11.7016$ $1.4\times 10^{-18}$ (✗)
MCAR $-2.3585$ $0.0209$ (✗) $-2.2937$ $0.0257$ (✗)
MNAR $-2.3697$ $0.0203$ (✗)
@@
MNAR $-2.0102$ $0.0479$ (✗) $0.1517$ $0.8798$
Time at Home (m) XO $-1.6511$ $0.1037$ $16.7191$ $7.4\times 10^{-24}$ (✗)
@@
MNAR $-1.0755$ $0.2854$ $-0.4124$ $0.6812$
Sleep Duration (s) XO $-3.0788$ $0.0029$ (✗) $22.2654$ $2.6\times 10^{-31}$ (✗)
@@
MNAR $-0.7592$ $0.4503$ $2.2641$ $0.0264$ (✗)
Time Walking (s) XO $-3.0425$ $0.0031$ (✗) $-4.1449$ $8.6\times 10^{-5}$ (✗)
@@
Location Clusters XO $-0.8132$ $0.4186$ $8.2048$ $6.9\times 10^{-12}$ (✗)
MCAR $-0.3650$ $0.7160$
MNAR $-0.7398$ $0.4616$ $1.5771$ $0.1189$
Total Steps XO $-0.1357$ $0.8924$ $5.2860$ $1.1\times 10^{-6}$ (✗)

### v3

@@ G.1 Signal Reconstruction and Imputation
Figure 11 presents reconstructed and imputed sample examples, where white shading indicates observed data, gray shading denotes originally missing data, and purple shading represents synthetically induced missingness. The remaining time steps (in this case, days) are fully visible to the model. When the original signal is obscured in observed intervals, it is due to one or more model reconstructions perfectly overlapping the true signal, demonstrating accurate recovery. As shown in Figures 11(a) and 11(b) all models perform well with binary variables.
Notably, the proposed VQ-TSFM and its variants E1 and E2 exhibit strong imputation capabilities even under high proportions of missingness, as evidenced by Figures 11(d), 11(f), 11(i), and 11(g). Whether the missing data spans large temporal segments (e.g., the first three-quarters of the sample in 11(f)), appears centrally (11(i)), or is intermittently distributed (11(g)), the models consistently maintain robust representations and plausible imputations. This performance generalizes across all variable types—continuous real-valued, continuous positive, count data, and binary—highlighting the versatility of the models across different data ranges and types. The implicit model—the simplest of the three proposed variants—exhibited performance on par with or better than that of E1 and E2. This indicates that, even without explicit conditioning on the missingness mask, VQ-TSFM effectively captures and reconstructs the underlying patterns of missing data.
@@ G.2 Embedding Usage Histograms
Sleep Start (s) XO 6.1860-6.1860 $3\times 10^{-8}$ (✗) 11.7016-11.7016 $1.4\times 10^{-18}$ (✗)
MCAR 2.3585-2.3585 $0.0209$ (✗) 2.2937-2.2937 $0.0257$ (✗)
MNAR 2.3697-2.3697 $0.0203$ (✗)
@@
MNAR 2.0102-2.0102 $0.0479$ (✗) $0.1517$ $0.8798$
Time at Home (m) XO 1.6511-1.6511 $0.1037$ $16.7191$ $7.4\times 10^{-24}$ (✗)
@@
MNAR 1.0755-1.0755 $0.2854$ 0.4124-0.4124 $0.6812$
Sleep Duration (s) XO 3.0788-3.0788 $0.0029$ (✗) $22.2654$ $2.6\times 10^{-31}$ (✗)
@@
MNAR 0.7592-0.7592 $0.4503$ $2.2641$ $0.0264$ (✗)
Time Walking (s) XO 3.0425-3.0425 $0.0031$ (✗) 4.1449-4.1449 $8.6\times 10^{-5}$ (✗)
@@
Location Clusters XO 0.8132-0.8132 $0.4186$ $8.2048$ $6.9\times 10^{-12}$ (✗)
MCAR 0.3650-0.3650 $0.7160$
MNAR 0.7398-0.7398 $0.4616$ $1.5771$ $0.1189$
Total Steps XO 0.1357-0.1357 $0.8924$ $5.2860$ $1.1\times 10^{-6}$ (✗)